### v2

@@ 1 Introduction
To address these issues, a third Veto station with vertical bars was installed during the Year End Technical Stop of 2023–2024 (see Figure 3). The acceptance of the Veto system was increased by lowering the whole system by . It now provides full coverage of the target area.
@@ 2 Design and construction
The Veto 3 module uses the same technology as Veto 1 and 2 [1].
While Veto 1 and 2 have horizontally disposed bars, the Veto 3 bars are aligned vertically (see Figure 4, top). This minimizes the probability that the readout dead time coincides with the passage of a second muon through the same bar. Although Veto 1 and Veto 2 are staggered, there remains a possibility that two muons could traverse the overlapping region. If the time interval between the two particles is sufficiently short, the signal from the second muon may be lost due to the time required to process the signal from the first. The vertical bars in Veto 3 provide an additional vertical segmentation, thereby reducing the overlapping region among the three bars. This minimizes the probability of two muons hitting the detector so close in space and time that no signal is left in any of the Veto planes.
The Veto 3 bars are long, extending by above and below the Veto 1 and 2 range (see Figure 3).
@@ 3 Performance with cosmic rays
The plot in Figure 6 shows the hit rate as a function of the hit position. It demonstrates that the hit rate does not depend upon the position along the bar. Indeed, the relative efficiency
@@
4 Installation
@@
After the completion of this civil engineering work, the upgraded Veto system was installed (see Figure 7).
5 Commissioning
The Veto 3 was commissioned using cosmic rays. The hit distribution per channel (see Figure 8) shows that the Veto plane was not damaged during transport and installation and that all channels were correctly taking data. The thresholds and gains from the TI18 configuration solve the noise and QDC saturation issue observed in the laboratory test (see Section 3).
@@ 6 Measured Veto inefficiency
The individual Veto plane inefficiencies are estimated using passing through muons by extrapolating reconstructed tracks in the SciFi and the DS to each plane and checking for signals (see [5]).
The results shown here are obtained using 51 of the delivered luminosity of which high-quality passing through muon events were selected.

### v3

@@ 1 Introduction
To improve the efficiency of the Veto system, a third Veto station with vertical bars was installed during the Year End Technical Stop of 2023–20244 (see Figure 3). Moreover, the whole Veto system was lowered by to provide full coverage of the target by all stations.
@@ 2 Design and construction
The Veto 3 module, made of scintillating bars read out by SiPMs, uses the same technology as Veto 1 and 2 [1].
While Veto 1 and 2 have horizontally disposed bars, the Veto 3 bars are aligned vertically (see Figure 4, top). The Veto 3 bars are long, extending by above and below the Veto 1 and 2 range (see Figure 3).
@@
The vertical disposition of bars in Veto 3 minimises the probability that the readout dead time coincides with the passage of a second muon through the same bar. Although Veto 1 and Veto 2 are staggered, there remains a possibility that two muons could traverse the overlapping region. If the time interval between the two particles is sufficiently short, the signal from the second muon may be lost due to the time required to process the signal from the first. The vertical bars in Veto 3 provide an additional vertical segmentation, thereby reducing the overlapping region among the three bars. This minimizes the probability of two muons hitting the detector so close in space and time that no signal is left in any of the Veto planes.
@@ 3 Performance with cosmic rays
The plot in Figure 6 shows the efficiency of each bar as well as the overall detector one as a function of the hit position. As expected, there is a minor and negligible effect due to the distance from the readout. Indeed, the relative efficiency
@@
4 Installation and commissioning
With the addition of the third plane, the overall thickness of the Veto system increased, requiring an adjustment of its position in both the upstream and vertical directions; to lower the system sufficiently to cover the target (see Figure 3), a trench was cut into the concrete floor.
@@ 4 Installation and commissioning
After completing this civil engineering work, the upgraded Veto system was installed. The final configuration is shown in Figure 7, which displays the new trench, the position of the Veto system relative to the rest of the detector, and the box housing its readout electronics.
The Veto 3 was commissioned using cosmic rays. The hit distribution per channel (see Figure 8) shows that the Veto plane was not damaged during transport and installation and that all channels were correctly taking data. The thresholds and gain settings adopted in the TI18 configuration effectively address the issues of electronic noise and QDC saturation that were identified during the laboratory testing phase (see Section 3). By carefully calibrating and tuning these parameters in TI18 conditions, the performance limitations encountered in the initial setup were mitigated, resulting in reduced noise and no QDC saturation.
@@ 5 Measured Veto inefficiency
The individual Veto plane inefficiencies are estimated using passing through muons by extrapolating reconstructed tracks in the SciFi and the DS to each plane. A plane is considered inefficient if it registers no hit where one is expected because of a passing track (see [6]).
The results shown here are obtained using of the delivered luminosity of which high-quality passing through muon events were selected.